In the same mouse species, a fourth unlinked gene (gene P/p) also affects fur color.
For mice that are either homozygous dominant (PP) or heterozygous (Pp), the organism’s fur color is dictated by the other three genes (A/a, B/b, and C/c).
For mice that are homozygous recessive (pp), large patches of the organism’s fur are white. This condition is called piebaldism.


In a cross between two mice that are heterozygous for agouti, black, color, and piebaldism, what is the probability that offspring will have solid black fur along withlarge patches of white fur?
Answer #1


In the same mouse species, a fourth unlinked gene (gene P/p) also affects fur color.
For mice that are either homozygous dominant (PP) or heterozygous (Pp), the organism’s fur color is dictated by the other three genes (A/a, B/b, andC/c).
For mice that are homozygous recessive (pp), large patches of the organism’s fur are white. This condition is called piebaldism.


In a cross between two mice that are heterozygous for agouti, black, color, and piebaldism,


what is the probability that offspring will have solid black fur along with large patches of white fur?

solid black fur--------probability ---aa---------1/4

black fur-------BB,bB--------------------3/4

COLORED FUR -probability --- Cc , CC----------3/4

PIE BALD -probability --- pp----------1/4

total probability of = 1/4 x 3/4 x 3/4x 1/4 = 9/256

answered by: zaib.u.nisa
Answer #2

thanks to answer #1

answered by: paypal